OpenStudy (anonymous):

HELP!!!!! A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Sn: 1^2 + 4^2 + 7^2 + . . . + (3n - 2)^2 = n(6n^2-3n-1)/2

OpenStudy (anonymous):

CAN YOU HELP ME...MY OLDER BRO MAKES ME DO HIS WORK SO I HAVE NO CLUE WHAT IM DOING

OpenStudy (cwrw238):

for S1 just plug n = 1 into the formula and see if it works out to 1^2 ( 1)

OpenStudy (cwrw238):

S1 = 1(6*1^2 - 3(1) - 1) / 2

OpenStudy (cwrw238):

what does that work out to?

OpenStudy (loser66):

that work out for S2, and S3 too. just plug n =2, and n =3 into the equation, you can see that it works for both them.

OpenStudy (anonymous):

17.5 ?

OpenStudy (cwrw238):

how did you get thatt?

OpenStudy (anonymous):

1(6*1^2 - 3(1) - 1) / 2 I DID THIS

OpenStudy (cwrw238):

ok that 1 * (6 - 3 -1) / 2 work out the brackets first = 1 * 2 / 2 = 2 / 2 = 1 which checks out

OpenStudy (cwrw238):

follow?

OpenStudy (anonymous):

YES

OpenStudy (cwrw238):

ok so do the same thing for S2 ( n = 2) and the result should equal the sum of 2 terms - that -s 1^2 + 4^2 which is 1 + 16 = 17

OpenStudy (anonymous):

SO WHAT WILL THE ANSWER BE @cwrw238 IM CONFUSED

OpenStudy (cwrw238):

plug in n = 2 into the formula Sn = n (6n^2 - 3n - 1) / 2 so S2 = 2(6*2^2 - 3*2 - 1 ) / 2 you just replace every n in the formula by 2 and work it out do brackets first what is 6* 2^2 - 3*2 - 1 equal to?

OpenStudy (anonymous):

17

OpenStudy (cwrw238):

do you know what the symbols mean what is 2^2 /

OpenStudy (cwrw238):

right 17

OpenStudy (anonymous):

DIVIDE

OpenStudy (cwrw238):

so we have 2 * 17 / 2

OpenStudy (anonymous):

17

OpenStudy (cwrw238):

= 34 / 2 = 17 thats the answer we require because S2 (th sum of the first 2 terms) = 1^2 + 4^2 = 17

OpenStudy (anonymous):

OK SO THEN WHAT YOU DID

OpenStudy (cwrw238):

yes work out the sum from the formula the check it against the sum you get by adding the number of terms so for S3 plug s=3 into the formula and that should work out to 1^2 + 4^2 + 7^2 = 1 + 16 + 49 = 66 if it does then the statement is true

OpenStudy (anonymous):

WHAT DO I SUM UP AGAIN?

OpenStudy (anonymous):

17 +17 +66?

OpenStudy (cwrw238):

no S3 means the sum of the first 3 terms that is 1^1 ^ 4^2 + 7^2 = 1 + 16 + 49 which equlas 66

OpenStudy (anonymous):

OH SO THE WHOLE ANSWER IS 66 @cwrw238 IM GOING TO THE 6TH GRADE SO I DONT KNOW WHAT THE HECK IM DOING

OpenStudy (cwrw238):

yes - now if you plug in n= 3 into the formula and work it out you should get 66 - this proves that he formula is true for S3.

OpenStudy (anonymous):

SO IS THIS WHAT MY ANSWER SHOULD LOOK LIKE S1 = 1(6*1^2 - 3(1) - 1) / 2=17 S2 = 1^2 + 4^2 = 17 S3 = 1^2 + 4^2 + 7^2 = 66 @cwrw238

OpenStudy (cwrw238):

S2 and S3 are correct but S1 = 1

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

S1 = 1(6*1^2 - 3(1) - 1) / 2=1 S2 = 1^2 + 4^2 = 17 S3 = 1^2 + 4^2 + 7^2 = 66

OpenStudy (cwrw238):

yea

OpenStudy (anonymous):

thank you so much @cwrw238

OpenStudy (cwrw238):

thats ok